Exercise 5-31 Methods of Estimating Costs: Account Analysis (LO 5-3) Records at Hals Accounting Services show the followingAdriana Corporation manufactures football equipment. In planning for next year, the managers want to understand the relation
Answer #1

a) Year 2 Cost

Direct material and supplies (47000*1.1) 51700
Direct labor (2600000*1.15) 2990000
Variable overhead (11.8*40000) 472000
Fixed overhead (730000*1.05) 766500
Total 4280200

b) Cost per billable hour

Cost per billable hour
Year 1 (3967000/50000) = 79.34
Year 2 (4280200/40000) = 107.01

2a) Variable cost per machine hour = Change in Cost/Change in machine hour = (109129-96969)/(1606-1288) = 38

Fixed cost = 109129-(1606*38) = 48101

2b) Total fixed cost = 48101+(1400*38) = 101301